Grade 12th passGeneral Physics

A particle is thrown vertically up such that distance travelled by it in 2nd and 9th second is seem maximum height upto which the particle Risers will be

Garima bhatt
8 Years agoGrade 12th pass

2 Answers

Shardool Pandit
8 Years ago
By Newton’s 2nd equation of motion,
Distance travelled in 2nd second = (2u - (10/2)*2*2) – (u - (10/2))
                                                    = u – 15
( here a = -10 as particle is moving up against gravity)
Distance travelled in 9th second = (9u + (10/2)*9*9) – (8u + (10/2)*8*8)
                                                    = u + 85
(here a = 10 because particle is moving down with gravity)
Now,  u – 15 = u + 85
          2u = 100
          u = 50 m/s
 
by Newton’s Third equation of motion,
v2 – u= 2as
While throwing up,
v = 0
u = 50 m/s
a = -10 m/s2
 
So, s = (0 – 2500) / (-20)
         = 125m ie Max height will be 125 metres.
mohammad uzair
6 Years ago
Here , we can directly use the formula for distance covered in nth second
i.e.,             sth= u+a/s(2n-1)
therefore...... For 2nd second ,
                                        S2nd=u+10/2(2×2-1) =u+ 15         …...........(1)
 similarly...........For 9th second
                                                 S9th=0+10/2(2×9-1) = 65            …............(2)
( in the later  case i have put u=0 because body is returning back after reaching to its highest point ,and at highest point velocity is zero)
Now equating 1 and 2 we will get 
                                           u= 50m/s
now usinng equation third of newton’s law of motion 
                                          v2= u2 + 2as
                                          (0)2 = (50)2 + 2 × 10 × s
                                            s= 125 m